myininaya (myininaya):

solve for :) :)^2+2:)+1=0

OpenStudy (saifoo.khan):

Math error! ^.^

OpenStudy (anonymous):

(:) + 1)^2 = 000000000000 :) = -111 +110

OpenStudy (anonymous):

Setting :) = x, Note: (+-) means plus or minus x=[-b(+-)sqrt(b^2-4ac)]/2a = [-2(+-)sqrt(4-4*4)]/4 = -2(+-)2i*sqrt(3)/4 = [1(+-)i*sqrt3)]/2 So x = 1/2 + i*sqrt(3)/2, or x = 1/2 - i*sqrt(3)/2. Since x = :), :) has the same values.